# Spurious-Aware Prototype Refinement for Reliable Out-of-Distribution Detection

### v1

@@ Appendix E Backbone Experiments
MLS $88.2_{{\color[rgb]{.5,.5,.5}\pm 2.8}}$ $98.1_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $95.1_{{\color[rgb]{.5,.5,.5}\pm 4.8}}$ $65.4_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $66.6_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $82.7$
@@
SHE $83.5_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $98.0_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $15.5_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $84.6_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $88.3_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $74.0$
@@
MSP $72.7_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $77.3_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $59.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $59.1_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $92.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $72.2$
@@
MLS $74.3_{{\color[rgb]{.5,.5,.5}\pm 4.2}}$ $77.7_{{\color[rgb]{.5,.5,.5}\pm 1.8}}$ $62.2_{{\color[rgb]{.5,.5,.5}\pm 4.0}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $92.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $72.9$
@@
RMDS $63.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $71.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $60.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $66.5_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $91.3_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $70.4$
@@
VIM $72.7_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $27.3_{{\color[rgb]{.5,.5,.5}\pm 3.0}}$ $39.2_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $86.5_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $51.5_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $55.4$
@@
ReAct $76.2_{{\color[rgb]{.5,.5,.5}\pm 4.6}}$ $77.7_{{\color[rgb]{.5,.5,.5}\pm 3.5}}$ $60.0_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $66.5_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $75.2$
@@
KNN $97.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $20.2_{{\color[rgb]{.5,.5,.5}\pm 5.4}}$ $82.3_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $88.0_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $49.9_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $67.7$
@@
MSP $82.3_{{\color[rgb]{.5,.5,.5}\pm 1.6}}$ $71.9_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $72.8_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $77.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.9$
@@
MLS $81.7_{{\color[rgb]{.5,.5,.5}\pm 2.8}}$ $73.1_{{\color[rgb]{.5,.5,.5}\pm 4.0}}$ $73.3_{{\color[rgb]{.5,.5,.5}\pm 7.8}}$ $78.2_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $95.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $80.4$
@@
KNN $92.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $30.6_{{\color[rgb]{.5,.5,.5}\pm 6.9}}$ $89.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $88.3_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $52.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $70.8$
@@
KLM $71.1_{{\color[rgb]{.5,.5,.5}\pm 6.9}}$ $77.8_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $59.2_{{\color[rgb]{.5,.5,.5}\pm 4.6}}$ $69.3_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $74.5$
@@
KLM $72.6_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 1.8}}$ $65.6_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $85.2_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.6$
@@
VIM $84.2_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $77.9_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $77.3_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $66.1_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $95.7_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $80.2$
@@
MDS $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $25.0_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $94.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $92.6_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $75.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $77.0$
@@
VIM $98.1_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $71.5_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $89.8_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $\textbf{88.9}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.6_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $89.2$
@@
Appendix F Analysis of Mixture of Prototypes
@@
As shown in Tables 34 and 35, SPROD-Default generally outperforms the two other variants when the amount of spurious correlation in the ID data is high (e.g., Waterbirds, CelebA, and UrbanCars); on other datasets, their performances are competitive. Proto-KMeans is slightly less effective in separating SP-OOD samples, likely due to its reliance on purely geometric clustering, whereas SPROD with classification-aware refinement slightly enhances the robustness when the spurious correlation is high.
@@
Appendix G Reproducibility and Resources
The code for our approach is available at the following anonymous GitHub repository: https://github.com/anonymouss-neurips2025/SPROD/. The cleaned AnimalsMetaCoCo dataset will be released upon acceptance of the paper and publication of the final version. In our benchmarking, we utilized components of the OpenOOD v1.5 [72, 37, 73] framework to obtain results for previously proposed OOD detection methods.

### v2

@@ Appendix G Backbone Experiments
MLS $88.2_{{\color[rgb]{.5,.5,.5}\pm 2.8}}$ $98.1_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $95.1_{{\color[rgb]{.5,.5,.5}\pm 4.8}}$ $5.4_{{\color[rgb]{.5,.5,.5}\pm 1.1}} 66$$6.6_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $82.7$
@@
SHE $83.5_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $98.0_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $15.5_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $4.6_{{\color[rgb]{.5,.5,.5}\pm 1.4}} 88$$8.3_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $74.0$
@@
MSP $2.7_{{\color[rgb]{.5,.5,.5}\pm 0.4}} 77$$7.3_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $59.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $59.1_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $92.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $72.2$
@@
MLS $4.3_{{\color[rgb]{.5,.5,.5}\pm 4.2}} 77$$7.7_{{\color[rgb]{.5,.5,.5}\pm 1.8}}$ $62.2_{{\color[rgb]{.5,.5,.5}\pm 4.0}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $92.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $72.9$
@@
RMDS $63.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $71.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $0.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 66$$6.5_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $91.3_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $70.4$
@@
VIM $72.7_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $27.3_{{\color[rgb]{.5,.5,.5}\pm 3.0}}$ $39.2_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $86.5_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $1.5_{{\color[rgb]{.5,.5,.5}\pm 1.2}} 55$.455.4
@@
ReAct $6.2_{{\color[rgb]{.5,.5,.5}\pm 4.6}} 77$$7.7_{{\color[rgb]{.5,.5,.5}\pm 3.5}}$ $0.0_{{\color[rgb]{.5,.5,.5}\pm 3.6}} 66$$6.5_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $75.2$
@@
KNN $97.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $20.2_{{\color[rgb]{.5,.5,.5}\pm 5.4}}$ $2.3_{{\color[rgb]{.5,.5,.5}\pm 0.4}} 88$$8.0_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $49.9_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $67.7$
@@
MSP $82.3_{{\color[rgb]{.5,.5,.5}\pm 1.6}}$ $71.9_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $2.8_{{\color[rgb]{.5,.5,.5}\pm 0.3}} 77$$7.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.9$
@@
MLS $81.7_{{\color[rgb]{.5,.5,.5}\pm 2.8}}$ $1_{{\color[rgb]{.5,.5,.5}\pm 4.0}} 73.3±7.$$3_{{\color[rgb]{.5,.5,.5}\pm 7.8}}$ $78.2_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $95.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $80.4$
@@
KNN $92.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $30.6_{{\color[rgb]{.5,.5,.5}\pm 6.9}}$ $9.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 88$$8.3_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $52.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $70.8$
@@
KLM $1.1_{{\color[rgb]{.5,.5,.5}\pm 6.9}} 77$$7.8_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $59.2_{{\color[rgb]{.5,.5,.5}\pm 4.6}}$ $69.3_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $74.5$
@@
KLM $2.6_{{\color[rgb]{.5,.5,.5}\pm 1.3}} 77$$7.4_{{\color[rgb]{.5,.5,.5}\pm 1.8}}$ $65.6_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $85.2_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.6$
@@
VIM $84.2_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $7.9_{{\color[rgb]{.5,.5,.5}\pm 1.5}} 77$$7.3_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $66.1_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $95.7_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $80.2$
@@
MDS $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $25.0_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $94.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $92.6_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $5.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 77$.077.0
@@
[…]
@@
Appendix H Analysis of Mixture of Prototypes
@@
As shown in Tables 35 and 36, SPROD-Default generally outperforms the two other variants when the amount of spurious correlation in the ID data is high (e.g., Waterbirds, CelebA, and UrbanCars); on other datasets, their performances are competitive. Proto-KMeans is slightly less effective in separating SP-OOD samples, likely due to its reliance on purely geometric clustering, whereas SPROD with classification-aware refinement slightly enhances the robustness when the spurious correlation is high.
@@
Appendix I Reproducibility and Resources
The data (cleaned AnimalsMetaCoCo dataset) and code for our approach are available at the following GitHub repository: https://github.com/ReihanehZohrabi/SPROD. In our benchmarking, we utilized components of the OpenOOD v1.5 [86, 37, 87] framework to obtain results for previously proposed OOD detection methods.